The area of a rectangle is 24 square meters.Find the length and width of the rectangle if its length is 2 meters greater than its width. Use an equation and the formula---->area of the rectangle=(width)(length).
Answer by jojo14344(1513)

Absolutely!
First, we know that Area(rect)=(Length)(Width)--- mark eqn. 1
And we know too that the Length is 2meters greater than width,
so L= W+2, right?---- mark eqn. 2 Then substitute this value of L in eqn 1,
A=(W+2)(W)
24= W^2+2W
W^2+2W-24=0
(W+6)(W-4)=0
It has 2 values,
W=-6, cannot be because there's no negative measurement for rect. right?
The other, W=4, perfect! This is the width.
Go back eqn 2 to get L= W+2=4+2
L=6 meters
In doubt? Go back eqn 1 to check,
A=LW
24=6*4
24sq m = 24sq m cool!
Thank you,
Jojo